# Identification of a Kalman Filter: Consistency of Local Solutions

### v2

@@ 1 Introduction
Identifying a model from measurements is an important task, especially for designing model-based controllers. To efficiently apply such algorithms, three requirements are central: an accurate predictive model, efficient online state estimation, and, sometimes, uncertainty quantification. Regarding the first requirement, it is often a mixture of prior knowledge from physics-based modeling and data-driven modeling. A popular approach for this task is parametric system identification using Prediction Error Methods (PEM) (Ljung, 2002) or Maximum Likelihood Estimation (MLE) (Åström, 1979; Simpson et al., 2023). Regarding the second requirement, for Linear Time-Invariant (LTI) systems, online state estimation is often performed using Kalman filters (Anderson and Moore, 1979). Such a filter requires knowledge of the process and measurement noise covariance matrices, which are often difficult to derive from the system’s physics. Several approaches exist to estimate them from data (Abbeel et al., 2005; Odelson et al., 2006), but if they are entirely unknown, it is often preferable to estimate the Kalman gain directly with PEM or MLE, possibly jointly with other parameters (Kuntz and Rawlings, 2025). While these methods have strong statistical guarantees, they require solving a generally non-convex optimization problem to global optimality for these guarantees to hold. This is a limitation because derivative-based optimization algorithms can only guarantee convergence to a local minimizer. A natural question arises: can we still provide statistical guarantees for local minimizers?
@@
This paper provides a positive answer to this question for the case of Kalman gain estimation using PEM. This follows from the fact that the optimization problem is asymptotically unimodal: as the amount of data goes to infinity, the limit of the objective function has a unique local (and therefore global) minimizer. This can be summarized as: non-global local minimizers are, asymptotically, not attributable to the estimation of the Kalman gain. We also propose some extensions of this result to more general cases, such as the joint identification of the innovation covariance matrix with MLE. However, one cannot provide guarantees for the completely general case because a poorly chosen parameterization can always lead to artificial local minima.
Literature Review.
@@
The asymptotic unimodality of PEM and MLE has been proven for a few specific classes of Single-Input Single-Output (SISO) systems. A summary of classic results is given in Ljung (1999, Section 10.5). A notable one is asymptotic unimodality for ARMA models (Åström and Söderström, 1974), which are black-box single-output autonomous LTI systems. Other results exist for SISO systems with specific structures or input design (Söderström, 1975; Goodwin et al., 2003; Zou and Heath, 2009; Eckhard et al., 2012). To the best of the authors’ knowledge, this paper provides the first asymptotic unimodality result for multiple-output systems (apart from the trivial case of linear regression), and the first time-domain analysis of this problem.
@@
In Section 2, we define the stochastic system of interest and the associated identification problem. In Section 3, we formulate PEM for this problem, where only the Kalman gain is estimated. We also propose stability-enforcing constraints, and recall the state-of-the-art result on the consistency of the global solutions. In Section 4, we state the main results of this paper: asymptotic unimodality and consistency of local minimizers. These theoretical results are illustrated with numerical examples in Section 5. Finally, in Section 6 we discuss possible extensions of these results to more general settings. We draw conclusions and discuss future research directions in Section 7.
@@
Throughout this paper, we use some common mathematical notations. We denote by $I$ the identity matrix of appropriate dimensions. The weighted norm associated with a positive-definite matrix $P$ is denoted by $\left\lVert x\right\rVert_{P}\coloneqq\sqrt{x^{\!\top}\!Px}$ and the matrix norm induced by the $L_{2}$-norm is denoted by $\left\lVert\cdot\right\rVert$. We use the matrix inequality notation $M_{1}{\,\preccurlyeq\,}M_{2}$ (resp. $M_{1}{\,\prec\,}M_{2}$) when the matrix $M_{2}-M_{1}$ is positive semi-definite (resp. positive-definite). The trace of a matrix $M$ is denoted by ${\mathrm{Tr}}{\left(M\right)}$. Finally, ${\mathrm{cl}}{\left({\mathcal{L}}\right)}$ denotes the closure of a set $\mathcal{L}$}.
@@ 2 Problem Statement
Here, $x_{k}\in{\mathbb{R}}^{{n}},u_{k}\in{\mathbb{R}}^{{p}},y_{k}\in{\mathbb{R}}^{{q}}$ and $e_{k}\in{\mathbb{R}}^{{q}}$ denote the state, the input, the output, and the innovation, respectively. The matrices $A\in{\mathbb{R}}^{{n}\times{n}}$, $B\in{\mathbb{R}}^{{n}\times{p}}$, and $C\in{\mathbb{R}}^{{q}\times{n}}$ as well as the initial state $x_{0}$ are known. The goal is to estimate the observer gain $L\in{\mathcal{L}}$ from the data $\{u_{k},y_{k}\}_{k=0,\dots,N}$, where ${\mathcal{L}}\subset{\mathbb{R}}^{{n}\times{q}}$ is some feasible set of gains.
@@
  5. e)
    The innovations $e_{k}$ are independent, zero-mean random variables with a constant positive covariance matrix and bounded fourth-order moments:
    ${\mathbb{E}}\left[e_{k}\right]=0,\quad{\mathbb{E}}\left[e_{k}e_{k}^{\top}\right]=S^{\star}{\,\succ\,}0,\quad{\mathbb{E}}\left[\left\lVert e_{k}\right\rVert^{4}\right]\leq c.$ (2)
the covariance matrix $S^{\star}{\,\succ\,}0$ is, in general, unknown, but we do not attempt to estimate it (apart from Section 6). Also, note that we do not need to assume that the experiment is open-loop, meaning that $u_{k}$ can be generated by a feedback controller that uses past outputs $y_{k-i}$.
@@ 3 The prediction error method
In our setting, the PEM (Ljung, 2002) leads to the following optimization problem:
@@
where the weighting matrix $W{\,\succ\,}0$ is some positive-definite matrix in ${\mathbb{R}}^{{q}\times{q}}$, and where the predicted states $\hat{x}_{k}(L)$ are obtained from the Kalman filter equations:
@@
A crucial observation is that the predicted states $\hat{x}_{k}(L)$ depend nonlinearly on the parameter $L$ because of the term $LC\hat{x}_{k}(L)$” in (4). In fact, this is what makes the optimization problem (3) non-convex in general.
Now we make an important assumption regarding the feasible set $\mathcal{L}$}.
@@
Note that, by a continuity argument, the inequality (5) also holds for $L\in{\mathrm{cl}}{\left({\mathcal{L}}\right)}$. Also, note that $\mathcal{L}$} is necessarily bounded because (5) implies $\left\lVert LC\right\rVert\leq\left\lVert A\right\rVert+\gamma$ and $C$ is full row-rank (cf. Assumption 1). Therefore, ${\mathrm{cl}}{\left({\mathcal{L}}\right)}$ is compact, which will be helpful later.
From a more practical point of view, there exist several ways to impose such a uniform stability constraint. For example, in Kuntz and Rawlings (2025), this is done via Linear Matrix Inequalities (LMI), and in Diehl et al. (2009), the spectral radius is approximated by a smooth function that involves some Lyapunov equation. Similarly, we propose to ensure (5) as follows:
@@
for some choice of $\alpha>0$. Note that this formulation is equivalent to imposing $\rho_{\alpha}(A-LC)\leq 1$ where $\rho_{\alpha}(\cdot)$ is the smooth spectral radius approximation used in Diehl et al. (2009). As remarked there, the constraint $P{\,\succcurlyeq\,}0$ is in fact never active (because $P{\,\succcurlyeq\,}I{\,\succ\,}0$ for any feasible point), so even though (6) is a nonlinear semi-definite program, it can be (almost) treated as an ordinary nonlinear program in practice.
Proposition 3 below draws a connection between (6) and Assumption 2, and the proof is provided in Appendix A.
@@
For any $\alpha>0$, the set ${\mathcal{L}}_{\alpha}$ defined below satisfies Assumption 2, and contains $L^{\star}$ for $\alpha$ small enough:
@@
Regarding the uniform convergence of $V_{N}(L)$, the proof relies on a lemma for stochastic dynamical systems presented in Ljung (1999, Theorem 2B.1) and repeated here in Appendix B.
@@ 4 Consistency of local solutions
The proof is provided in Appendix C. Note that the value of $V_{N}(L)$ also converges uniformly to $\bar{V}(L)$ over $L\in{\mathcal{L}}$, but we do not need this assertion here.
@@
Since the pair $[A,C]$ is observable (cf. Assumption 1.a), the pair $A-\hat{L}C,C]$ is also observable (as a consequence of the Hautus lemma). Hence, (18) implies that D=0{D}=0. We continue as follows:
@@
Let, for all $N\in{\mathbb{N}}$, $\hat{L}_{N}\in{\mathcal{L}}$ be stationary points of $V_{N}(\cdot)$. Then $\hat{L}_{N}$ is a strongly consistent estimate of $L^{\star}$, i.e.,
@@
The results of this section rely on Assumptions 1 and 2. Even if a feasible set $\mathcal{L}$} satisfying Assumption 2 is not explicitly used, Theorem 6 still holds provided that the sequence $\{\hat{L}_{N}\}_{N\in{\mathbb{N}}}$ satisfies the uniform stability condition (5). We can even go further, and claim that if the objective value $V_{N}(\hat{L}_{N})$ remains bounded, then the condition (5) will almost surely be satisfied for $N$ large enough. This claim is unfortunately not proven here; we leave it for future work.
@@ 5 Numerical Examples
A two-state example.
We next study a system with two states and a single measurement. The underlying continuous-time dynamics describe a particle subject to linear friction and a random piecewise-constant force:
$\ddot{q}(t)$ $=-\mu\dot{q}(t)+f_{k},\quad t\in[k\Delta t,(k+1)\Delta t),$ (22)
where $q(t)$ is the position of the particle at time $t$, $\mu>0$ is the friction coefficient, and $f_{k}\sim{\mathcal{N}}(0,\,\sigma_{f}^{2})$ is a random force that remains constant over each sampling period of length $\Delta t>0$. The measurements take the form $y_{k}=q(k\Delta t)+v_{k}$, where $v_{k}\sim{\mathcal{N}}(0,\,\sigma_{v}^{2})$ is the measurement noise. This system is discretized analytically into a discrete-time LTI model with $x_{k}=[q(k\Delta t),\dot{q}(k\Delta t)]^{\!\top}\!$, which we can put in the innovation form (1) after computing the true Kalman gain $L^{\star}$ from the Discrete Algebraic Riccati Equation (DARE) (Anderson and Moore, 1979). To estimate the gain, we solve the constrained PEM problem (6) with the Lyapunov-based stability constraint with some constant $\alpha\in(0,1)$. In practice, we optimize over $L$ only by eliminating $P$ via (6b). Derivatives are computed with CasADi, and an interior-point method (Nocedal and Wright, 2006) is implementedfootnotemark: with line-search and a Gauss-Newton Hessian approximation.
@@
Figure 2: Optimization iterates and solutions for the two-state example with $\alpha=0.02$, and the model parameters $\mu=0.1$, $\Delta t=0.1$, $\sigma_{f}=10$, and $\sigma_{v}=1$.
@@
A three-state position-acceleration example.
Finally, we consider a slightly more realistic problem with three states and two measurements. The continuous-time dynamics are again given by (22), but the external force $f_{k}$ now evolves according to a first-order stochastic model, and both position and acceleration are measured:
$y_{k}=\begin{bmatrix}\ddot{q}(k\Delta t)+v^{{\mathrm{acc.}}}_{k}\\ q(k\Delta t)+v^{{\mathrm{pos.}}}_{k}\end{bmatrix}\qquad f_{k+1}$ $=a_{f}f_{k}+w_{k}.$ (23)
Measurement noises are Gaussian, while the process noise $w_{k}$ follows a mixture distribution: $w_{k}\sim{\mathcal{N}}(0,\,\sigma_{w}^{2})$ with probability $p>0$, and $w_{k}=0$ with probability 1p1-p. Note that $w_{k}$ still has zero mean, finite fourth-order moments, and a positive variance ${\mathbb{E}}\left[w_{k}^{2}\right]=p\sigma_{w}^{2}$.
@@
Figure 3: Estimation error for the three-state example with two measurements, for different realizations and different values of $N$. The model parameters are as in Figure 2, with in addition: ${\mathrm{Cov}}\left[v^{{\mathrm{acc.}}}_{k}\right]=1$, ${\mathrm{Cov}}\left[v^{{\mathrm{pos.}}}_{k}\right]=2$, $p=0.1$, $\sigma_{w}^{2}=10$, $a_{f}=0.9$.
The results show a clear decrease in the estimation error as $N$ grows, despite the non-Gaussian process noise. This empirical behavior is consistent with the strong consistency of local minimizers established in Theorem 6. We also observe a convergence speed of order $\mathcal{O}(\frac{1}{\sqrt{N}})$ which is consistent with the law of large numbers that was used in the proof.
Finally, as empirical evidence of the asymptotic unimodality of the cost, we generate $50$ feasible initial guesses using Kalman filters with random matrices $Q$ and $R$ and observe that all initializations lead to the same solution (up to tolerance) when $N=100$.

### v3

@@ 1 Introduction
Identifying a model from measurements is an important task, especially for designing model-based controllers. To efficiently apply such algorithms, three requirements are central: an accurate predictive model, efficient online state estimation, and, sometimes, uncertainty quantification. Regarding the first requirement, it is often a mixture of prior knowledge from physics-based modeling and data-driven modeling. A popular approach for this task is parametric system identification using Prediction Error Methods (PEM) (Ljung, 2002) or Maximum Likelihood Estimation (MLE) (Åström, 1979; Simpson et al., 2023). Regarding the second requirement, for Linear Time-Invariant (LTI) systems, online state estimation is often performed using Kalman filters (Anderson and Moore, 1979). Such a filter requires knowledge of the process and measurement noise covariance matrices, which are often difficult to derive from the system’s physics. Several approaches exist to estimate them from data (Abbeel et al., 2005; Odelson et al., 2006), but if they are entirely unknown, it is often preferable to estimate the Kalman gain directly with PEM or MLE, possibly jointly with other parameters (Kuntz and Rawlings, 2025). This motivates analyzing PEM for Kalman gain estimation. While this method has strong statistical guarantees, they require solving a generally non-convex optimization problem to global optimality for these guarantees to hold. This is a limitation because derivative-based optimization algorithms can only guarantee convergence to a local minimizer. A natural question arises: can we still provide statistical guarantees for local minimizers?
@@
This paper provides a positive answer to this question for the case of Kalman gain estimation using PEM. This follows from the fact that the optimization problem is asymptotically unimodal: as the amount of data goes to infinity, the limit of the objective function has a unique local (and therefore global) minimizer. Thus, asymptotically, non-global local minimizers are not attributable to the estimation of the Kalman gain. While we do not analyze the finite-sample case, in our experience, a few samples are usually sufficient to observe the unimodality, as illustrated by numerical examples. We also propose some extensions of this result to more general cases, such as the joint identification of the innovation covariance matrix with MLE. However, one cannot provide guarantees for the completely general case because a poorly chosen parameterization can always lead to artificial local minima.
@@
The asymptotic unimodality of PEM and MLE has been proven for a few specific classes of Single-Input Single-Output (SISO) systems. A summary of classic results is given in Ljung (1999, Section 10.5). A notable one is asymptotic unimodality for ARMA models (Åström and Söderström, 1974), which are black-box single-output autonomous LTI systems. Other results exist for SISO systems with specific structures or input design (Söderström, 1975; Goodwin et al., 2003; Zou and Heath, 2009; Eckhard et al., 2012). To the best of our knowledge, this paper provides the first asymptotic unimodality result for multiple-output systems (apart from the trivial case of linear regression), and the first time-domain analysis of this problem.
@@
In Section 2, we define the stochastic system and identification problem of interest. In Section 3, we formulate PEM for this problem, with stability-enforcing constraints, and recall the state-of-the-art result on the consistency of the global solutions. In Section 4, we state the main results of this paper: asymptotic unimodality and consistency of local minimizers. These theoretical results are illustrated with numerical examples in Section 5. Finally, in Section 6, we discuss possible extensions of these results to more general settings. Conclusions and future research directions are discussed in Section 7.
@@
Throughout this paper, we denote by $I$ the identity matrix of appropriate dimensions. We use $\left\lVert x\right\rVert_{P}\coloneqq\sqrt{x^{\!\top}\!Px}$ to denote the weighted norm associated with $P{\,\succ\,}0$, and omit the subscript when $P=I$. The trace of a matrix $M$ is denoted by ${\mathrm{Tr}}{\left(M\right)}$. Finally, ${\mathrm{cl}}{\left({\mathcal{L}}\right)}$ denotes the closure of a set $\mathcal{L}$}.
@@ 2 Problem Statement
Here, $x_{k}\in{\mathbb{R}}^{{n}},u_{k}\in{\mathbb{R}}^{{p}},y_{k}\in{\mathbb{R}}^{{q}}$ and $e_{k}\in{\mathbb{R}}^{{q}}$ denote the state, the input, the output, and the innovation, respectively. The goal is to estimate the observer gain $L\in{\mathcal{L}}\subset{\mathbb{R}}^{{n}\times{q}}$ from the data $\{u_{k},y_{k}\}_{k=0,\dots,N}$.
The key motivation for this innovation form is that, under mild assumptions, LTI systems with process and measurement noise can be equivalently represented in this form, with $L$ being the Kalman gain.
@@
  6. f)
    The innovations $e_{k}$ are zero-mean independent random variables such that:
    ${\mathbb{E}}\left[e_{k}e_{k}^{\top}\right]=S^{\star}{\,\succ\,}0,\quad\text{and}\quad\sup_{k\in{\mathbb{N}}}{\mathbb{E}}\left[\left\lVert e_{k}\right\rVert^{4}\right]<\infty.$ (2)
The covariance matrix $S^{\star}{\,\succ\,}0$ is, in general, unknown, but we do not attempt to estimate it (apart from Section 6). Note that assumption e) is not restrictive as the Kalman gain $L^{\star}$ is always stabilizing the system when $(A,C)$ is observable.
@@ 3 The prediction error method
In the presented settings, the PEM (Ljung, 2002) leads to the following optimization problem:
@@
where the weighting matrix $W\in{\mathbb{R}}^{{q}\times{q}}$ is some positive-definite matrix, and the predicted states $\hat{x}_{k}(L)$ are obtained from the Kalman filter equations:
@@
A crucial observation is that the predicted states $\hat{x}_{k}(L)$ depend nonlinearly on the parameter $L$ because of the term $LC\hat{x}_{k}(L)$” in (4). In fact, this is what makes the optimization problem (3) non-convex in general. Now we make an important assumption regarding the feasible set $\mathcal{L}$}.
@@
Importantly, $\mathcal{L}$} is necessarily bounded because (5) implies $\left\lVert LC\right\rVert\leq\left\lVert A\right\rVert+\gamma$ and $C$ is full row-rank (cf. Assumption 1). Therefore, ${\mathrm{cl}}{\left({\mathcal{L}}\right)}$ is compact, which will be helpful later. By a continuity argument, (5) also holds for all $L\in{\mathrm{cl}}{\left({\mathcal{L}}\right)}$.
From a more practical point of view, there exist several ways to impose such a uniform stability constraint. For example, in Kuntz and Rawlings (2025), this is done via Linear Matrix Inequalities (LMI). Similarly, we propose to ensure (5) using some Lyapunov equation constraint, as it is done in Diehl et al. (2009):
@@
for some $\alpha>0$. In the our experience, this method is quite robust against the choice of parameter $\alpha$: it only needs to be small enough for $L^{\star}$ to be feasible in (6). We choose $\alpha=0.02$ in the numerical examples.
The constraints (6b)-(6d) are equivalent to $\rho_{\alpha}(A-LC)\leq 1$ where $\rho_{\alpha}(\cdot)$ is the smooth spectral radius approximation used in Diehl et al. (2009). As noted there, the constraint $P{\,\succcurlyeq\,}0$ is in fact never active (because $P{\,\succcurlyeq\,}I{\,\succ\,}0$ for any feasible point), so even though (6) is a nonlinear semi-definite program, it can be (almost) treated as an ordinary nonlinear program in practice.
Proposition 3 below draws a connection between (6) and Assumption 2. A proof is provided in the Appendix.
@@
For any $\alpha>0$, the set ${\mathcal{L}}_{\alpha}$ defined below satisfies Assumption 2:
@@
Regarding the uniform convergence of $V_{N}(L)$, the proof relies on a lemma for stochastic dynamical systems presented in Ljung (1999, Theorem 2B.1) and repeated in the Appendix of the present paper.
@@ 4 Consistency of local solutions
The proof is provided in the Appendix. Note that the value of $V_{N}(L)$ also converges uniformly to $\bar{V}(L)$ over $L\in{\mathcal{L}}$, but we do not need this assertion here.
@@
Since the pair $[A,C]$ is observable (cf. Assumption 1.a), the pair $A-\hat{L}C,C]$ is also observable (this is a consequence of the Hautus lemma Hautus (1969)). Hence, (18) implies that D=0{D}=0. We continue as follows:
@@
Let, for all $N\in{\mathbb{N}}$, $\hat{L}_{N}\in{\mathcal{L}}$ be stationary points of $V_{N}(\cdot)$.
Then, $\hat{L}_{N}$ is a strongly consistent estimate of $L^{\star}$:
@@
The results of this section rely on Assumptions 1 and 2. Even if a feasible set $\mathcal{L}$} satisfying Assumption 2 is not explicitly used, Theorem 6 still holds provided that the sequence $\{\hat{L}_{N}\}_{N\in{\mathbb{N}}}$ satisfies the uniform stability condition (5), for some $\gamma>0$ and $\lambda\in(0,1)$. This is the case whenever $\hat{L}_{N}$ are bounded, and $\rho(A-\hat{L}_{N}C)$ is bounded away from $1$. As a conjecture, we also claim (without a proof) that this condition holds whenever the objective value $V_{N}(\hat{L}_{N})$ remains bounded.
@@ 5 Numerical Examples
A two-state example with scalar measurements.
We next study a system with two states and scalar measurements. The underlying continuous-time dynamics describe a particle subject to linear friction and a random piecewise-constant force:
$\ddot{q}(t)$ $=-\mu\dot{q}(t)+w_{k},\quad t\in[k\Delta t,(k+1)\Delta t),$ (22)
where $q(t)$ is the position of the particle at time $t$, $\mu>0$ is the friction coefficient, and $w_{k}\sim{\mathcal{N}}(0,\,\sigma_{w}^{2})$ is a random force that remains constant over each sampling period of length $\Delta t>0$. The measurements take the form $y_{k}=q(k\Delta t)+v_{k}$, where $v_{k}\sim{\mathcal{N}}(0,\,\sigma_{v}^{2})$ is the measurement noise. This system is discretized analytically into a discrete-time LTI model with $x_{k}=[q(k\Delta t),\dot{q}(k\Delta t)]^{\!\top}\!$, which we can put in the innovation form (1) after computing the true Kalman gain $L^{\star}$ from the Discrete Algebraic Riccati Equation (DARE) (Anderson and Moore, 1979). To estimate the gain, we solve the constrained PEM problem (6) with the Lyapunov-based stability constraint with some constant $\alpha\in(0,1)$. In practice, we optimize over $L$ only by eliminating $P$ via (6b). Derivatives are computed with CasADi, and an interior-point method (Nocedal and Wright, 2006) is implementedfootnotemark: with line-search and a Gauss-Newton Hessian approximation.
@@
Figure 2: Optimization iterates and solutions for the two-state example with $\alpha=0.02$, and the model parameters $\mu=0.1$, $\Delta t=0.1$, $\sigma_{w}=10$, and $\sigma_{v}=1$.
@@
A two-state example with two measurements.
Finally, we consider a slightly more realistic problem: the same model as (22) but where both position and acceleration are being measured:
$y_{k}=\begin{bmatrix}q(k\Delta t)+v_{k}&\qquad\ddot{q}(k\Delta t)+v^{\prime}_{k}\end{bmatrix}^{\top}$. (23)
Measurement noises are still Gaussian: $v_{k}\sim{\mathcal{N}}(0,\,\sigma_{v}^{2})$ and $v^{\prime}_{k}\sim{\mathcal{N}}(0,\,\sigma_{v^{\prime}}^{2})$. The process noise $w_{k}$ follows a mixture distribution: $w_{k}\sim{\mathcal{N}}(0,\,\sigma_{w}^{2})$ with probability $p>0$, and $w_{k}=0$ with probability 1p1-p. Note that $w_{k}$ still has zero mean, finite fourth-order moments, and a positive variance ${\mathbb{E}}\left[w_{k}^{2}\right]=p\sigma_{w}^{2}$.
@@
Figure 3: Estimation error in the example with two measurements, for 1010 realizations. The model parameters are as in Figure 2, and the noise parameters are $\sigma_{v}=2$, $\sigma_{v^{\prime}}=1$, $\sigma_{w}=1$ and $p=0.1$.
The results show a clear decrease in the estimation error as $N$ grows, despite the non-Gaussian process noise. This empirical behavior is consistent with the strong consistency of local minimizers established in Theorem 6. We also observe a convergence speed of order $\mathcal{O}(1/\sqrt{N})$ which is consistent with the law of large numbers that was used in the proof.
Finally, for $N=100$, we solve the problem starting from $50$ different initial guesses generated using Kalman filters with random matrices $Q$ and $R$. We observe that all distances between the solutions are less than $10^{-4}$, which is the numerical tolerance of the solver. This suggests that all initializations lead to the same minimizer, which provides empirical evidence of the asymptotic unimodality of the cost.